OpenStudy (anonymous):

f(t)=200(5-(9/2+t)) find the average rate of change of f(t) during the first year, use MVT? Any help would be appreciated

OpenStudy (anonymous):

there arent any intervals given which makes this problem harder

OpenStudy (anonymous):

MVT MEANS ?

OpenStudy (anonymous):

this looks weird

OpenStudy (anonymous):

Mean value theorem

OpenStudy (anonymous):

it is really \[f(t)=200(5-(9/2+t))\]?

OpenStudy (anonymous):

the average rate of change refers to the mean value theorem but I dont have any intervals so would it be DNE?

OpenStudy (anonymous):

yeah it is unfortunately...

OpenStudy (anonymous):

or maybe \[f(t)=200(5-\frac{9}{2+t})\]?

OpenStudy (anonymous):

the one that you just put up is correct

OpenStudy (anonymous):

it says "first year" so assuming \(t\) is given in years since you start counting, compute \[f(1)-f(0)\]

OpenStudy (anonymous):

ahh so itd be [0,1]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

no need to divide, since length of interval is 1

OpenStudy (anonymous):

awesome thanks

OpenStudy (anonymous):

yw